HIGH RESOLUTION INFRARED SPECTRA OF LARGER HELIUM CLUSTERS: He_{N} - OCS WITH N UP TO 70

\underline{A.R.W. McKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

Recently it has been possible to study the onset of superfluid behavior in small (N ∼ 10) [1] and nanoscale (N ∼ 1000)

[2] helium clusters by observing the spectrum of a probe molecule like OCS, whose vibrational and rotational motions are a sensitive indication of the local helium environment. Our previous IR and microwave spectroscopy [1] of small He_{N} - OCS clusters extended up to N = 8. With CO or CO_{2} as the probe [3] it was possible to approach N = 20. Using a new apparatus with a partially skimmed pulsed supersonic jet expansion, it has now been possible to resolve and assign distinct IR spectra of He_{N} - OCS for virtually every single N-value from 1 to over 70. The observed lines remain sharp (<0.001 cm^{−1}) at least up to this cluster size. Analysis of these vibration-rotation spectra, in the 2062 cm^{−1} region of the OCS ν_{1} fundamental band, show that the cluster B-values exhibit an unexpected oscillatory behavior which experimentally marks the completion of the second and third solvation shells of helium around the OCS. At N = 70, the cluster properties are already fairly close to those of large nanodroplets, meaning that we have bridged much of the gap between the individual molecule and bulk matter worlds on a one-N-at-a-time basis.

[1] J. Tang, Y. Xu, A.R.W. McKellar, and W. Jäger, Science 297, 2030 (2002).

[2] S. Grebenev, M. Hartmann, M. Havenith, B. Sartakov, J.P. Toennies, and A.F. Vilesov, J. Chem. Phys. 112, 4485 (2000).

[3] J. Tang and A.R.W. McKellar, J. Chem. Phys. 119, 754 (2003); 121, 181 (2004).